# mycolorRGB219, 48, 122

### v1

@@ 2.5 Characterizing curvature and morphology
The scaling argument described above suggests the dominance of sheet-like structures. Motivated by this prediction, we measured the surface curvature of the active/passive interfaces. We compute the values of the two principal curvatures $\kappa_{1}(\mathbf{s})$ and $\kappa_{2}(\mathbf{s})$, where $\mathbf{s}$ is a point on the interface $\mathcal{S}$. We choose $\kappa_{1}<\kappa_{2},$ everywhere on the phase-separated interface (Fig. 5A) (Materials and Methods). We show representative visualizations of the normalized Gaussian curvature $w_{a}^{2}\kappa_{1}(\mathbf{s})\kappa_{2}(\mathbf{s})$ for $\mathbf{s}\in\mathcal{S},$ where $w_{a}$ is to the width of the active phase, as defined in the previous section. In both experiments and simulations, the surface generated by active phase separation is dominated by regions of normalized Gaussian curvature close to zero (Fig. 5 B,C). We contrast this with the transient morphology obtained in passive phase separation from the Cahn-Hilliard equation, which is dominated by saddle-like surfaces with negative Gaussian curvature  (?, ?, ?, ?, ?) (Fig. 5D).
From the principal curvature values, we compute the joint probability distributions $\rho(\kappa_{1},\kappa_{2})$. Different regions in the $\kappa_{1}$-$\kappa_{2}$ plane indicate different surfaces (Fig. 5E). In both experiments and simulations, the joint distribution is peaked at small values of both $\kappa_{1}$ and $\kappa_{2},$ which corresponds to sheet-like domains (Fig. 5F, G). The vertical and horizontal branches centered around zero-curvature corresponds to cylinder-like surfaces, which are associated with the sheet edges. In comparison, the probability distribution obtained from the Cahn-Hilliard equation is characterized by large positive-negative pairs of principal curvatures correspond to saddle-like surfaces (Fig. 5H). The curvature analysis supports the finding that the connected active domain is dominated by sheet-like structures.
@@
Figure 5: Comparing the interface morphology in passive and active bicontinuous structures. (A) The two principal curvatures $\kappa_{1}$ and $\kappa_{2}$ are calculated everywhere on the interface. (B-D) 3D visualization of the normalized Gaussian curvature $\kappa_{1}\kappa_{2}w_{a}^{2}$ of the interface from experiment (B), simulations of the active/passive mixture(C) and simulations of the passive Cahn-Hilliard equation (D). The color indicates the value of the Gaussian curvature. (E-H) Distribution $\rho(\kappa_{1},\kappa_{2})$ of normalized principal curvatures $\kappa_{1}w_{a}$ and $\kappa_{2}w_{a}$. Different values of $\kappa_{1}$ and $\kappa_{2}$ correspond to different geometries (E). $\rho$ peaks at small values of $\kappa_{1}$ and $\kappa_{2}$ in experiments (F) and simulations (G), indicating sheet-like interfaces. $\rho$ peaks at $\kappa_{1}\ll 0,\kappa_{2}\gg 0$ in simulations of the Cahn-Hilliard equation (H), indicating a saddle-like interface. The experimental sample (B,F) contains 92 nM KSA, 1.8% PEO, 1.8% dextran, and $\phi_{a}=0.38$. Simulation parameters (C,G): $|\alpha|=1.30,\gamma=1.0,$ and $\phi_{a}=0.40$.
@@ 3 Discussion
Activity transforms the coarsening domains of passive three-dimensional phase separation into a steady state bicontinuous structure that continuously reconfigures in time, while maintaining its morphology. The bicontinuous morphology is observed in a dynamical steady-state across a wide range of volume fractions, while its width is controlled by activity and surface tension. Active bicontinuous structure is composed of sheet-like domains, in contrast to saddle-like interfaces observed in passive phase separation   (?, ?, ?, ?, ?).
There is a stark contrast between how bicontinuous structures form in passive and active fluids   (?, ?). In both cases, passive restorative forces such as interfacial tension drive the system to a state of minimal surface area. In a passive system, established strategies to preserve bicontinuity in steady state therefore either engineer the system’s chemistry, like tailoring molecular interactions in block copolymers, or kinetically arrest the structure by solidifying one phase or the interface itself (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). In contrast, active fluid maintains a dynamical, bicontinuous steady state over its entire lifetime, which is determined by the available chemical fuel.
@@
The effect of activity on phase separation has been explored both experimentally and theoretically. In particular, extensive theoretical work has been carried out in scalar models of active/passive mixtures, where the mixture is described entirely in terms of the conserved concentration field, coupled to flow  (?, ?, ?). In this limit, active stresses are driven by gradients of the concentration field and are therefore localized at the interface. These models exhibit arrested coarsening driven by the competition between active stress and interfacial tension. In a different limit studied here, isotropic or nematic liquid crystal and their associated active stresses are present throughout the bulk of the active fluid  (?, ?, ?, ?, ?, ?). An important interaction for the formation of bicontinuous structures is active anchoring at the interface  (?), where the active units align along the interfaces to generate autonomous flows that stretch the active phase into sheet-like domains (Fig. 5). This elongation continuously connects separate droplets into a network and dynamically balances the collapsing effect of surface tension. Whether bicontinuous structures can be obtained in three-dimensional scalar models, perhaps as a result of active renormalization of surface tension, remains, to our knowledge, an open question.
@@
The structures in active phase separation are determined by the competition between active stress and surface tension. We have shown that this competition controls the width of the active sheets (Fig. 4). Extreme values of activity, however, break bicontinuity resulting in distinct behaviors. On the one hand, when activity is very small compared to surface tension, active flows cannot significantly elongate the droplets, but only induce small interfacial deformations. In this case, one of the two phases irreversibly turns into droplets through a Plateau-Rayleigh instability  (?, ?). On the other hand, when activity is very high, the associated energy input is large enough to overcome the effect of the thermodynamic chemical potential, suppressing the phase separation itself and generating a uniform mixture (?, ?).
@@
Biological and biomimetic systems are known for their flexibility and the capability to self-heal. This stems from their use of continuous energy input, which allows them to maintain dynamical steady states, adapt to damage, and resist degenerative processes such as coarsening. Our work demonstrates how this principle can be applied to synthetic systems to generate controllable structures not achievable in equilibrium.
@@ 4.2 Experimental Design
The phase separation system and the MT-based active system were based on previous protocols (?, ?), with a decreased concentration of coexisting polymers to reduce surface tension. PEO (100 kDa, Sigma-Aldrich) and dextran (450 to 650 kDa, Sigma-Aldrich) were added into M2B buffer (80 mM K-pipes, 2 mM MgCl2, 1 mM EGTA, pH 6.8). We varied their initial concentrations to three values during our experiment: 1.8% PEO and 1.8% dextran, 1.9% PEO and 1.9% dextran and 2.0% PEO and 2.0% dextran, which controlled the capillary length of the sample (Fig. S1). For visualization, we added a very small amount (<<0.1% w/w) of amino-dextran (Fina Biosolutions) fluorescently labeled with Alexa-Fluor 488 NHS Ester (Invitrogen).
Tubulin was purified and labeled as described in previous work  (?). A minor tubulin fraction (3%) was conjugated to AZDye 647 NHS Ester (Fluoroprobes) using a succinimidyl ester linkage. Labeled and unlabeled tubulin monomers were co-polymerized into microtubules (MTs) at 37 C using 0.6 mM guanylyl-(alpha,beta)-methylene-diphosphonate (GMPCPP, Gena Biosciences) and 1.2 mM Dithiothreitol (DTT, Gena Biosciences) in M2B buffer (?). The MTs were sonicated for 60 seconds in a Branson 1800 sonicator to reduce the bundle rigidity . The final concentration of MTs was 0.67 mg/mL. Kinesin-401 protein fused to a biotin-carboxyl carrier domain was expressed and purified using established protocols (?). This protein was subsequently combined with tetrameric streptavidin in an M2B buffer with 5 mM DTT, forming a kinesin-streptavidin cluster (KSA).The final concentrations of KSA in the samples varied from 42-183 nM.
@@
The dextran phase had a slightly larger density than the PEO phase, causing a sedimentation in the z-z direction due to gravity over tens of minutes. This would finally lead to two bulk phase separated phases, with the active phase perforated with passive droplets at the bottom of the chamber. To minimize this effect, after the samples were activated, they were placed on a rotator (Benchmark Scientific H2020 Roto-Therm) and rotated at 24 rpm at room temperature for 1.5 hours. This ensured that the samples were in steady state without sedimentation (Fig. S4). The samples were then immediately transferred to the microscope and imaged consecutively for 5 minutes (Fig. S3).
@@ 4.3 Data Analysis
Funding:
This work was primarily supported by the US Department of Energy, Office of Basic Energy Sciences under award number DE-SC0019733. Development and optimization of the two-phase system of active liquid–liquid phase separation was supported by NSF-ISS-2224350. OS and MT acknowledge support of NSF-MRSEC DMR 2308708. We also acknowledge the use of the biosynthesis facility supported by NSF-MRSEC-2011846 and NRI-MCDB Microscopy Facility supported by the NIH Shared Instrumentation Grant 1S10OD019969-01A1. Use was made of computational facilities purchased with funds from the NSF(CNS-1725797) and administered by the Center for Scientific Computing (CSC). The CSC is supported by the California NanoSystems Institute and the Materials Research Science and Engineering Center (NSF DMR 2308708).
Author contributions:
P.G., L.Z., Z.D., and M.C.M designed research; P.G., and L.Z. performed research; P.G., L.Z., and M.T. analyzed data; P.G., L.Z., M.T., O.A.S, Z.D., and M.C.M. wrote the paper.
Competing interests:
The authors declare that they have no competing interests.
@@
Figs. S1 to S3
Tables S1 to S4
References (7-0)
Movie S1
Data S1
@@
Dynamical bicontinuous networks from 3D active phase separation
@@
Figure S4: Bicontinuous structure reached steady state. 3D visualizations of the dextran phase (red) at 4 different time points (A-D). The network structure reached a steady state at $t=40$ min and remained similar over 2 hours (B-D) together with the active width $w$ (E). The sample initially contain 92 nM KSA, 1.8% PEO and 1.8% dextran and is recombined to obtain active fraction $\phi_{a}=0.28$.
Figure S5: Bicontinuity for increasing activity The dependence of active/passive domain connectedness fractions on the activity. For large activities, the bicontinuous state for the entire range of volume fractions explored numerically i.e. the spinodal region of the passively phase separating system.
Figure S6: Visualization of active phase (red) morphology over varying $\phi_{a}$. Samples contain $\phi_{a}=$0.16 (A), 0.29 (B), 0.38 (C), 0.60 (D), corresponding to data points in Fig. 2 D. All samples initially contain 183 nM KSA, 1.8% PEO, and 1.8% dextran and are recombined to obtain various $\phi_{a}$.
@@
Three-dimensional visualization of active phase (red) forming a dynamic network in experiment. The movie starts at $1.5$ h after the sample was prepared. The sample initially contains 1.91.9% PEO, 1.91.9% dextran, $183$ nM KSA.

### v2

@@ 2.5 Characterizing curvature and morphology
The scaling argument described above suggests the dominance of sheet-like structures. To investigate further, we quantify the surface curvature in both experiment and simulation. We compute the values of the two principal curvatures $\kappa_{1}(\mathbf{s})$ and $\kappa_{2}(\mathbf{s})$, where $\mathbf{s}$ is a point on the interface $\mathcal{S}$. We define $\kappa_{1}(s)<\kappa_{2}(s)$ for each $s$ (Fig. 5A) (Methods 4.3). We visualize the interface structure with its normalized Gaussian curvature $w_{a}^{2}\kappa_{1}(\mathbf{s})\kappa_{2}(\mathbf{s})$ for $\mathbf{s}\in\mathcal{S},$ where $w_{a}$ is the width of the active phase, as defined in the previous section. In both experiments and simulations, the interface generated by active phase separation is dominated by regions of normalized Gaussian curvature close to zero (Fig. 5 B,C). We contrast this with the transient morphology obtained in passive phase separation from the Cahn-Hilliard equation, which is dominated by saddle-like surfaces with negative Gaussian curvature  (?, ?, ?) (Fig. 5D).
@@
Figure 5: Comparing the interface morphology in passive and active bicontinuous structures. (A) The two principal curvatures $\kappa_{1}$ and $\kappa_{2}$ are calculated everywhere on the interface. (B-D) 3D visualization of the normalized Gaussian curvature $\kappa_{1}\kappa_{2}w_{a}^{2}$ of the interface from experiment (B), simulations of the active/passive mixture(C) and simulations of the passive Cahn-Hilliard equation (D). The color indicates the value of the Gaussian curvature. (E-H) Distribution $\rho(\kappa_{1},\kappa_{2})$ of normalized principal curvatures $\kappa_{1}w_{a}$ and $\kappa_{2}w_{a}$. Different values of $\kappa_{1}$ and $\kappa_{2}$ correspond to different geometries (E). $\rho$ peaks at small values of $\kappa_{1}$ and $\kappa_{2}$ in experiments (F) and simulations (G), indicating sheet-like interfaces. $\rho$ peaks at $\kappa_{1}\ll 0,\kappa_{2}\gg 0$ in simulations of the Cahn-Hilliard equation (H), indicating a saddle-like interface. The experimental sample (B,F) contains 92 nM KSA, 1.8% PEO, 1.8% dextran, and $\phi_{a}=0.37$. Simulation parameters (C,G): $|\alpha|=1.30,\gamma=1.0,$ and $\phi_{a}=0.40$.
From the principal curvature values, we compute the joint probability distributions $\rho(\kappa_{1},\kappa_{2})$. Different regions in the $\kappa_{1}$-$\kappa_{2}$ plane indicate different structures (Fig. 5 E). In both experiments and simulations, the joint distribution is peaked at small values of both $\kappa_{1}$ and $\kappa_{2},$ which corresponds to sheet-like domains (Fig. 5 F,G). The vertical and horizontal branches centered around zero-curvature corresponds to cylinder-like surfaces, which are associated with the sheet edges. In comparison, the probability distribution obtained from the Cahn-Hilliard equation is characterized by large positive-negative pairs of principal curvatures correspond to saddle-like surfaces (Fig. 5H). The curvature analysis supports the finding that the connected active domain is dominated by sheet-like structures.
@@ 3 Discussion
We showed how activity transforms the coarsening domains of passive three-dimensional phase separation into a continuously reconfiguring bicontinuous structure that maintains its morphology. The bicontinuous morphology is observed across a wide range of volume fractions, while its width is controlled by activity and surface tension. The active bicontinuous structure is composed of sheet-like domains, in contrast to saddle-like interfaces observed in passive phase separation  (?, ?, ?, ?, ?).
@@
Extensive previous work examined the effect on activity on phase separation has been explored from multiple perspectives. In one regime, also studied here, active isotropic or nematic liquid crystals coexist with a passive fluid  (?, ?, ?, ?, ?, ?). In such systems, active stresses are present throughout the bulk of the active fluid. System-spanning active networks over a wide range of volume fractions have been studied in two-dimensional simulations  (?). Building on these findings, we use simulations and experiments to demonstrate that bicontinuity is a ubiquitous feature of three-dimensional active phase separation. An important ingredient required for the formation of bicontinuous structures is active anchoring  (?), where the active units align along the interfaces to generate autonomous flows that stretch the active phase into sheet-like domains. This elongation continuously connects separate droplets into a network and dynamically counters the collapsing effect of surface tension. Theoretical work also explored a different regime, described by scalar models of active/passive mixtures, where the mixture is described in terms of the conserved concentration field, coupled to flow  (?, ?, ?). In this regime, active stresses are driven by gradients of the concentration field that are localized at the interface. Whether bicontinuous structures can be obtained in three-dimensional scalar models, perhaps as a result of active renormalization of surface tension, remains, to our knowledge, an open question.
@@
The structures in active phase separation are determined by the competition between active stress and surface tension. We have shown that this competition generates bicontinuous morphology and controls the width of the active sheets. The bicontinuity, however, is only observed over a wide but well-defined range of activities. On the one hand, when activity is very small compared to surface tension, active flows cannot elongate the droplets, but only induce small interfacial deformations. In this case, one phase irreversibly turns into droplets through a Plateau-Rayleigh instability  (?). On the other hand, for very high activities, the energy input overcomes the effect of the thermodynamic chemical potential, suppressing the phase separation itself and generating a uniform mixture (?, ?). Furthermore, bicontinuous structures are only observed in a finite range of active fraction $\phi_{a}$ (Fig. 2). At low fractions of the active fluid elongated droplets cannot merge. In this regime, the system transitions from droplets to a uniform phase with decreasing surface tension, without an intermediate bicontinuous state  (?). Finally, gravity causes a non-uniform distribution of the two phases, generating extreme densities $\phi_{a}$ which do not allow bicontinuity  (?). Our work provides specific techniques and parameters to overcome these difficulties and achieve a controllable bicontinuous fluid.
Bicontinuous structures also arise in passive phase separation  (?). However, without active forces, passive restorative interfacial tension drives the system to a state with a minimal surface area. The characteristic length of conventional coarsening dynamics grows with a power law $w\propto t^{\nu_{p}}$, where $\nu_{p}$ varies between 1/3 and 1  (?, ?, ?). Therefore, bicontinuous structures in passive systems are transient. In contrast, the active stresses maintain a dynamical, bicontinuous steady state throughout the entire active fluid lifetime.
@@
Biological systems are adaptable and self-healing. Using continuous energy input, they maintain dynamical steady states, adapt to damage, and resist degenerative processes. Our work demonstrates how using continuous energy input can generate controllable and highly dynamical biomimetic structures and morphologies that are not easily achievable in equilibrium.
@@ 4.2 Experimental Design
The experimental system of active-passive phase separation is based on previous protocols  (?, ?), with a decreased concentration of coexisting polymers to reduce surface tension. PEO (100 kDa, Sigma-Aldrich) and dextran (450 to 650 kDa, Sigma-Aldrich) were added into M2B buffer (80 mM K-pipes, 2 mM MgCl2, 1 mM EGTA, pH 6.8). We studied three initial concentrations: 1.8% PEO and 1.8% dextran, 1.9% PEO and 1.9% dextran and 2.0% PEO and 2.0% dextran. Decreasing polymer concentration decreased the capillary length and surface tension (Fig. S1). For visualization, we added (<<0.1% w/w) of amino-dextran (Fina Biosolutions) fluorescently labeled with Alexa-Fluor 488 NHS Ester (Invitrogen).
Tubulin was purified and labeled as described previously  (?). A minor tubulin fraction (3%) was conjugated to AZDye 647 NHS Ester (Fluoroprobes) using a succinimidyl ester linkage. Labeled and unlabeled tubulin monomers were co-polymerized into microtubules (MTs) at 37 C using 0.6 mM guanylyl-(alpha,beta)-methylene-diphosphonate (GMPCPP, Gena Biosciences) and 1.2 mM Dithiothreitol (DTT, Gena Biosciences) in M2B buffer (?). The MTs were sonicated for 60 seconds in a Branson 1800 sonicator to reduce filament length, which generated softer bundles that fragment more easily. The final MT concentration was 0.67 mg/mL. Kinesin-401 protein fused to a biotin-carboxyl carrier domain was expressed and purified using established protocols  (?). Kinesin-401 was combined with tetrameric streptavidin in an M2B buffer with 5 mM DTT, forming a kinesin-streptavidin cluster (KSA). The final concentrations of KSA in the samples varied from 42-183 nM.
@@
The denser dextran phase sediments in the z-z direction over tens of minutes. This would generate two bulk phase separated phases, with the active phase perforated with passive droplets at the bottom of the chamber. To minimize this effect, samples were placed on a rotator (Benchmark Scientific H2020 Roto-Therm) and rotated at 24 rpm at room temperature for 1.5 hours. This reduced sedimentation (Fig. S4). For imaging, the samples were transferred to the microscope and imaged for a maximum of 5 minutes (Fig. S3).
@@ 4.3 Data Analysis
Figs. S1 to S6
Movies S1 to S2
@@
Bicontinuity in active phase separation
@@
Figure S4: Bicontinuous structure reached steady state. 3D visualizations of the dextran phase (red) at 4 different time points (A-D). The network structure reached a steady state at $t=40$ min and remained similar over 2 hours (B-D) together with the active width $w_{a}$ (E). The sample initially contain 92 nM KSA, 1.8% PEO and 1.8% dextran and is recombined to obtain active fraction $\phi_{a}=0.28$.
Figure S5: Bicontinuity for increasing activity. The dependence of active/passive domain connectedness fractions on the activity. For large activities, the bicontinuous state for the entire range of volume fractions explored numerically i.e. the spinodal region of the passively phase separating system.
Figure S6: Visualization of active phase (red) morphology over varying $\phi_{a}$. Samples contain $\phi_{a}=$0.16 (A), 0.29 (B), 0.37 (C), 0.60 (D), corresponding to data points in Fig. 2D. All samples initially contain 183 nM KSA, 1.8% PEO, and 1.8% dextran and are recombined to obtain various $\phi_{a}$.
@@
Three-dimensional visualization of active phases (red) forming dynamic networks in experiment at different KSA concentrations. The movie starts at $1.5$ h after all sample were prepared. The samples initially contain 1.81.8% PEO, 1.81.8% dextran, 4218342-183 nM KSA.